RC Meeting, Tuesday, October 05 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  1. No major showstoppers from the accelerator side yesterday.
  2. Today, Ops is doing beam studies started at 09:00 and expected to conclude at 13:00. During this time, there will be no beam into the Hall and any pending work can get done. Beam recovery is expected to start at 13:00 (this is NOT the time when Halls should expect beam)

Hall C Status

  1. Oct 04 Swing (yesterday):
    • At 13:24, there was an HMS Dipole Interlock alarm --> HMS Dipole Interlock. Contacted Amy Comer, but nothing could be done remotely. She contacted Joe Beaufait and he went to the Hall with RADCON to check the problem. Joe and Ellen managed to temporarily fix the issue by 17:00 --> HMS Dipole Fixed
    • At ~19:57, SHMS Q1 went into interlock --> SHMS Q1 Interlock and I contacted Ellen Becker (the new magnet expert on-call). Ellen went into the Hall and fixed. She also went ahead and fixed the minor issue with the HMS Dipole ITrue readback --> SHMS Q1 Fixed
  2. Oct 05 Owl/Day (today):
    • At 07:34, I was informed about SHMS Q1 Interlock re-occurrence. Contacted Ellen and she will need to go into the Hall. Shift crew has requested a full survey of the Hall so that Ellen can have more time to look at issue more in-depth in both the HMS as well as SHMS magnets as interlocks have been more frequent. Accelerator will also be down during this time (starting at 09:00 am) for maintenance until 13:00, when beam recovery will start.

Current Run Plan

  1. After beam studies concludes today at 13:00 and Ops restores beam into Hall C, Dave Mack would like to do a BCM Calibration (any time after 15:00 hrs). We should let Hall B know beforehand, as these will be invasive to them.
  2. Continue taking production data on LH2 target with 45 uA beam at new kinematics (SHMS: 8.035 GeV/c, 5.505 deg | HMS: -1.026 GeV/c, 47.07 deg)
    • As of 07:27 today (Owl Shift), we have accumulated: 1975 pions (19.75 %) (thanks Richard)

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  2. HMS Dipole NMR GUI displays Itrue = 0, even though it was set to a specified current --> HMS Dipole Readback . This issue was addressed by Ellen Becker, during the Controlled Access that was done to fix the SHMS Q1 Interlock yesterday. --> HMS Dipole Readback Fixed ( CLEARED )
  3. Radiation rope set by RADCON near HMS will certainly cause problems if the HMS is rotated, as the rope passes right through the HMS center.
    • I have notified the Crew Chief who will let RADCON know about this as I could not contact RADCON. Stephen Kay checked, and the the rope is actually part of a spool attached to the HMS so that when it rotates, the spool extends. ( CLEARED )

Special Requests

  1. Contact Ellen Becker if there is an opportunity to access the hall (~ 4 hours). This is for diagnosing the SHMS power supply. ( CLEARED )
    • Given the recurrent magnet interlocks in the Oct 05 Owl Shift, and that the Accelerator is down for maintenance, Ellen will take this opportunity to diagnose the SHMS, and given the time, the HMS magnet power supplies to try and resolve the cause of the interlocks. As of 13:00, Ellen has replaced the faulty power supply with a new brand, and she suggests that the new brand of power supply is used, as it is suspected that the issue is with the previous brand. See log entry --> SHMS Q1 Interlock Fixed